1.in 2009 Usain bolt set a new world record time of 9.58 seconds for the mens 100 metres. What was his average speed for the race

2.kenyan athlete Daniel Komen holds the world record for the mens 3000 metres. if his average speed for the race was 6.81 metres per second, what is his world record time? give your answer in minutes and seconds.
1
Expert's answer
2019-04-29T10:16:52-0400

1.


"v=\frac{100}{9.58}=10.4\frac{m}{s}"

2.


"t=\frac{3000}{6.81}=440.53 s=7 \ min \ 20.5 s"